OpenStudy (anonymous):

Factor completely. 20x3 + 12x2 + 5x + 3

OpenStudy (anonymous):

20x3 + 12x2 + 5x + 3 take out common from the first two terms 4x^2(5x+3)+1(5x+3) (4x^2+1)(5x+3)

OpenStudy (anonymous):

We can do this: \[20x^3 + 12x^2 + 5x + 3=4(5x^3 + 3x^2)+ 5x + 3=4x^2(5x+3)+5x+3\] How can we continue now?

OpenStudy (anonymous):

@steff06 hope you get it

OpenStudy (anonymous):

@steff06 This was useful? Please tell us if you have any doubt.